A 25-ft ladder is placed 8 ft from the base of the house. How
high up the side of the house can the ladder reach? Round
your answer to the nearest foot.

Answer by nerdybill(7384)

A 25-ft ladder is placed 8 ft from the base of the house. How
high up the side of the house can the ladder reach? Round
your answer to the nearest foot.
.
Let x = reach up side of house
then
x^2 + 8^2 = 25^2
x^2 = 25^2 - 8^2
x^2 = 625 - 64
x^2 = 561
x = 23.68
rounding:
x = 24 feet